OpenStudy (anonymous):

You have the numbers 1–24 written on slips of paper. If you choose one slip at random, what is the probability that you will not select a number which is divisible by 3?

OpenStudy (unklerhaukus):

\[ {\{3,6,9,12,15,18,21,24\} \over \{1,2,3,4,5,6,7,8,9,10,11,12,13,14,15,16,17,18,19,20,21,22,23,24 \}} \] \[=1/3\]

OpenStudy (unklerhaukus):

that is 1/3 probability you select a number divisible by three, the probability that the number is not divisible by three is the total probability - 1/3 \[=1-1/3=2/3\]